OpenStudy (yananeedshelp):

Which statements describe one of the transformations performed on f(x)=x^2 to create g(x)=3(x+5)^2-2? A. A translation of 5 units to the left. B. A translation of 2 units to the left. C. A vertical stretch with a scale factor of 3. D. A vertical stretch with a scale factor of 1/3.

OpenStudy (mathmate):

Hint: From f(x)=x^2 (or any other function) g(x)=f(x-h)+k, where h,k \(\in R\), i.e. h & k are real numbers. means translation to the right if h>0 translation to the left if h<0 translation upwards if k>0 translation downwards if k<0. Example: f(x)=x^2+3x g(x)=f(x-3)+7 ...... h=3, k=7, ....translating 3 units right, 7 units up. =(x-3)^2+3(x-3) + 7 = x^2-3x+7

OpenStudy (yananeedshelp):

Thanks!